OpenStudy (anonymous):

2(4x-5)+2=10-4(6x-45)

OpenStudy (texaschic101):

do you know how to distribute through the parenthesis ?

OpenStudy (anonymous):

Yeah lol Im a math wiz when it came to algebra xD I wanted to know if you know it

OpenStudy (texaschic101):

take the number before the parenthesis and multiply it by everything in the parenthesis. 2(4x - 5) would be 8x - 10 ( I multiplied 2 by 4x and then multiplied 2 by -5. That gets rid of the parenthesis. Now lets do this problem.. 2(4x - 5) + 2 = 10 - 4(6x - 45) 8x - 10 + 2 = 10 - 24x + 180 8x - 8 = -24x + 190 8x + 24x = 190 + 8 32x = 198 x = 198/32 x = 6.1875

OpenStudy (texaschic101):

or x = 6 6/32 which reduces to 6 3/16

OpenStudy (anonymous):

I know.. Im not stupid xD

OpenStudy (texaschic101):

I didn't say you were :)

OpenStudy (anonymous):

Yeah but your acting like it >:O

OpenStudy (texaschic101):

why do you post questions that you know the answer to ??

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.